# Projected Augmented Waves (PAW) inspired mixed basis sets for small molecules

### v1

@@ Abstract
The success behind many pseudopotential methods such as Projected Augmented Waves (PAW) and Phillips-Kleinman pseudopotential methods is that these methods are all electron methods in disguise. For the Phillips Kleinman and PAW methods we show that there is an all electron formulation. In the all electron formulation there are regular low wavevector wavefunctions and several, specially chosen, high wavevector wavefunctions that are specialized to the atomic environment. Using this as inspiration here we propose a new, PAW pseudopotential inspired, basis set method for small molecules, where we use the recently introduced LO basis set is paired with a Gaussian basis for a hybrid basis for small molecules.
@@ I Introduction
It is of paramount importance to make progress on the basis set problem for Density Functional Theory (DFT) for small molecules. Small molecules larger then diatomic: carbon dioxide, water, ammonia, methane, ethyne, ethene, ethane to name a few have a variety of uses in industry and engineering [2, 1]. However there has been limited progress in optimizing basis sets for small molecule DFT calculations similarly to basis sets that have paved breakthroughs in solid state DFT [3, 4, 5]. Here inspired by pseudopotential methods in DFT we propose one such hybrid basis set. Pseudopotentials at their core map a problem into another problem. The exact all electron wavefunction for the system is mapped onto a smooth wavefunction that may well be represented by a small number of plane waves. Two of the pre-eminent examples of pseudopotentials are the Phillips Kleinman and Projected Augmented Waves (PAW) pseudopotentials. Here we show that these pseudopotential methods are all electron methods in disguise where the all electron problem includes smooth plane wave wave functions and specialized wave functions for orbitals near atomic nuclei. Inspired by this observation, and the extensive numerical success of the PAW method [3, 4] we propose new basis sets for small molecules, where localized LO or lo [6, 5, 7, 1] basis wave functions are augmented with Gaussian wave functions (where Gaussians have been extensively used for the problem and various configuration integrals for them are well known [8, 2, 3]), in a similar manner to Phillips-Kleinman and PAW all electron basis sets We propose that this new hybrid basis set is efficient for DFT and Hartree-Fock (HF) calculations of small molecules.
@@ II Many pseudopotential methods are all electron methods in disguise (motivation)
In this section we reformulate PAW and the Phillips-Kleinman method as all electron methods where there are plane waves and special wave functions that are chosen to represent states near the core. We show the two methods to be equivalent in their all electron formulation.
@@ II.1 Phillips-Kleinman as an all electron method in disguise
We now recall the Orthogonalized Plane Wave (OPW) wave functions used for the Phillips-Kleinman method. We know that
@@ II.2.1 PAW review
We would like to study highly oscillatory wave functions near the nucleus of an atom using plane waves. In order to obtain a low cutoff of say ~30 Rydberg in energy and as such a tolerable basis set we must have that the wavefunction we represent using plane waves is very smooth. The key idea of PAW is then to introduce:
@@
where $\hat{\mathcal{T}}$ is some linear transformation (specified below) and $\left|\psi_{n}\right\rangle$ is the exact Khon Sham wave functions (solutions of the KS equations) while $\left|\tilde{\psi}_{n}\right\rangle$ are the wave functions we approximate efficiently using plane waves. Now suppose there is a set of augmentation spheres and we wish for $\hat{\mathcal{T}}$ to be the identity outside the augmentation spheres, so we write:
@@
Here $\mu$ are the atomic spheres. Now we want the wavefunction $\left|\tilde{\psi}\right\rangle$ to be smooth inside the sphere and we know that the exact KS wavefunction is close enough to the single atom, atomic wavefunction $\left|\phi_{i}\right\rangle$. As such if we choose $S_{\mathbf{\mu}}$ to be
@@ II.2.2 Reformulation of PAW as an all electron method
for small number of plane waves here $\left|\psi_{n}\right\rangle$ are all electron wave functions. Now because $\left|\tilde{\phi}_{i}\right\rangle$ is very smooth we have that
@@
III GLOW and PLOW basis sets
@@
III.1 Glow and GLOW for small molecules
Here G in Glo and GLO stands for Gaussians while lo and LO for localized orbitals. We consider the lo basis set:
$\Phi_{lo}^{lm}=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (14)
@@ III.1 Glow and GLOW for small molecules
$\left[-\frac{d^{2}}{dr^{2}}+\frac{l\left(l+1\right)}{r^{2}}+\bar{V}_{KS}\left(r\right)\right]ru_{l\mu}^{E}\left(r\right)=Eru_{l\mu}^{E}\left(r\right)$ (15)
and $\bar{V}_{KS}\left(r\right)$ is the spherically average KS potential or HF single particle potential. Alternatively we can use LO basis sets
$\Phi_{LO}^{lm}$ $=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$
@@
$\phi_{l,m,\sigma}^{\mu}=\frac{1}{\mathcal{N}}\exp\left(-\frac{\left|\mathbf{r-r}_{\mu}\right|^{2}}{2\sigma}\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (17)
to the molecular basis set as well as polynomials times $\phi_{l,m,\sigma}^{\mu}$, these help make configuration integrals cheap. Here $\mathcal{N}$ is a normalization constant. This is a basis set for many small molecules.
III.2 Plow and PLOW basis sets for crystalline solids
@@
Here P in Plo and PLO stands for plane wave while lo and LO stands for localized orbitals. We start with lo basis wave functions
$\Phi_{lo}=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (18)
where $S_{l}^{\mu}$ is chosen to make the wavefunction continuous everywhere and LO basis wave functions:
$\Phi_{LO}$ $=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$
$+B_{l}\dot{u}_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (19)
where $B_{l}^{\mu}$ and $S_{l}^{\mu}$ are chosen to make the wavefunction continuous and continuously differentiable everywhere. Higher derivative terms are also possible [6, 11, 10]. We can also add plane wave
@@
to the solid state basis set, these help make configuration integrals cheap. This is a basis set for many crystalline solids.
@@ IV Conclusions
In this work - inspired by the fact that many pseudopotential methods are all electron methods in disguise with slowly oscillating wave functions and special wave functions to account for the environment of the nucleus - we proposed new basis sets for small molecules and crystalline solids. In this basis set we use the LO or lo wave functions introduced recently in [6] (although regular LO or lo basis wave functions will also do [5, 7, 1]) combined either with Gaussians (for molecules) or plane waves (for solids) to obtain a total basis for the system. This would allow for initial calculations for many molecules as the four center calculation is done for Gaussians and two center density density is done for very generic functions [2, 9, 8]. This should help open small molecules for theoretical DFT or HF like calculations and explorations.

### v3

@@ Abstract
The success behind many pseudopotential methods such as Projected Augmented Waves (PAW) and Phillips-Kleinman pseudopotential methods is that these methods are all electron methods in disguise. For the Phillips Kleinman and PAW methods we show that there is an all electron formulation. In the all electron formulation there are regular low wavevector basis wave functions and several, specially chosen, high wavevector basis wave functions that are specialized to the atomic environment. Using this as inspiration here we propose a new, PAW method inspired, basis set for small molecules, where we use the recently introduced LO or lo (Localized orbitals) basis set is paired with a Gaussian basis, Gaussian type orbitals (GTO), for a hybrid basis for small molecules. Contracted Gaussian functions are also possible (CGFs). We also show how to extend this idea to Slater basis set, Slater type orbitals (STO), instead of GTO orbitals combined with LO or lo orbitals.
@@ I Introduction
It is of paramount importance to make further progress on the basis set problem for Density Functional Theory (DFT) for small molecules. Small molecules larger then diatomic: carbon dioxide, water, ammonia, methane, ethyne, ethene, ethane to name a few have a variety of uses in industry and engineering [2, 1]. There has been as much progress in optimizing basis sets for small molecule DFT calculations, similarly to basis sets that have paved breakthroughs in solid state DFT [3, 4, 5, 9, 11, 12, 17, 14, 13, 8, 7, 6, 1, 10, 15, 16]. Much progress has been made using ideas based on Linear Combinations of Atomic Orbitals (LCAO) [19, 2, 18] where a good basis for the entire molecule consists of linear combinations of basis functions for the individual atoms. There are many candidate LCAO basis sets:
  • Gaussian basis sets (GTO) based on cartesian Gaussian Type Orbitals (GTO) which are of the form of Gaussians multiplied by polynomials in the cartesian co-ordinates [20, 18, 21, 2, 22, 23]. This is advantageous as four center integrals are efficiently done for this basis set.
  • Slater basis sets where Slater Type Orbitals (STOs) are used in the LCAO method. These are spherical harmonics multiplied by exponentially decaying radial wave functions. It is possible to have multiple wave functions per angular momentum channel - so called multi-zeta basis sets. STO’s have advantages over GTOs as they have the correct exponentially decaying behavior at infinity and the right Kato cusp at the origin [18, 2], however four center integrals are much harder.
  • Contracted basis sets (contracted Gaussian Functions (CGF)) where multiple Gaussians are combined into one basis element used to mimic STO basis elements while retaining the ability to efficiently study four centre integrals [20, 18, 21, 2].
Furthermore plane waves, outside the LCAO concept, are often used to study periodically arranged arrays of molecules, with the distance between molecules on the order of 10 $nm$ so that interactions between the molecules may be well neglected. The most common electronic structure method for this situation (plane waves) is Projected Augmented Waves (PAW) [24, 25, 3, 4].
Here we focus on the LCAO method. Inspired by pseudopotential methods in DFT we propose one such (LCAO) hybrid basis set. Pseudopotentials at their core map a problem into another problem. The exact all electron wavefunction for the system is mapped onto a smooth wavefunction that may well be represented by a small number of plane waves, that is it is on its face not all electron. Two of the pre-eminent examples of pseudopotentials are the Phillips Kleinman and Projected Augmented Waves (PAW) pseudopotentials. Here we show that these pseudopotential methods are all electron methods in disguise, where the all electron problem includes smooth plane wave wave functions and specialized wave functions for orbitals near atomic nuclei. Inspired by this observation, and the extensive numerical success of the PAW method [3, 4] we propose new basis sets for small molecules, where localized LO or lo [26, 5, 6, 1] basis wave functions are augmented with Gaussian wave functions - GTOs - (where Gaussians have been extensively used for the problem and various configuration integrals for them are well known [18, 2, 3]) or STO’s for enhanced accuracy. This is done in a manner similar to Phillips-Kleinman and PAW all electron basis sets. We propose that this new hybrid basis set is efficient for DFT and Hartree-Fock (HF) calculations of small molecules.
@@ II Many pseudopotential methods are all electron methods in disguise (motivation)
In this section we reformulate PAW and the Phillips-Kleinman method as all electron methods. We show there are plane waves and special wave functions that are chosen to represent states near the core. We show the two methods to be equivalent in their all electron formulation depending on the choice of the localized wave functions.
@@ II.1 Phillips-Kleinman as an all electron method in disguise
We now recall the Orthogonalized Plane Wave (OPW) wave functions used for the Phillips-Kleinman method. We know that:
@@ II.2.1 PAW review
We would like to study highly oscillatory wave functions near the nucleus of an atom using plane waves. In order to obtain a low cutoff of say ~50 Rydberg in energy and as such a tolerable basis set we must have that the wavefunction we represent using plane waves is very smooth. The key idea of PAW is then to introduce:
@@
where $\hat{\mathcal{T}}$ is some linear transformation (specified below) and $\left|\psi_{n}\right\rangle$ is the exact Khon Sham wave functions (solutions of the KS equations) while $\left|\tilde{\psi}_{n}\right\rangle$ are the wave functions we approximate efficiently using plane waves. Now suppose there is a set of augmentation spheres and we wish for $\hat{\mathcal{T}}$ to be the identity outside the augmentation spheres (centered around nuclei) , so we write:
@@
Here $\mu$ are the atomic augmentation spheres. Now we want the wavefunction $\left|\tilde{\psi}\right\rangle$ to be smooth inside the sphere and we know that the exact KS wavefunction is close enough to the single atom, atomic wavefunction $\left|\phi_{i}\right\rangle$. As such if we choose $S_{\mathbf{\mu}}$ to be
@@ II.2.2 Reformulation of PAW as an all electron method
to significant accuracy, for small number of plane waves here $\left|\psi_{n}\right\rangle$ are all electron wave functions. Now because $\left|\tilde{\phi}_{i}\right\rangle$ is very smooth we have that
@@
III GLO, SLO and PLO basis sets
@@
III.1 Glo and GLO for small molecules
Here G in Glo and GLO stands for Gaussians while lo and LO for localized orbitals. We consider the lo basis set [26]:
$\Phi_{lo}^{\mu Elm}=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (14)
@@ III.1 Glo and GLO for small molecules
$\left[-\frac{d^{2}}{dr^{2}}+\frac{l\left(l+1\right)}{r^{2}}+\bar{V}_{KS}\left(r\right)\right]ru_{l\mu}^{E_{l}}\left(r\right)=E_{l}ru_{l\mu}^{E_{l}}\left(r\right)$ (15)
and $\bar{V}_{KS}\left(r\right)$ is the spherically average KS potential or HF single particle potential. Furthermore $E_{l}$ is chosen at the binding energy of the single atom system (other options are possible). Alternatively we can use LO basis sets [26]:
$\Phi_{LO}^{\mu Elm}$ $=u_{l}^{E_{l}}\left(\left|\mathbf{r}-\mathbf{r}_{\mu}\right|\right)\Theta\left(S_{l}^{\mu}-\left|\mathbf{r-r}_{\mu}\right|\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$
@@
$G_{l,m,\sigma}^{\mu}=\frac{1}{\mathcal{N}_{l}}\exp\left(-\frac{\left|\mathbf{r-r}_{\mu}\right|^{2}}{2\sigma}\right)\left|\mathbf{r-r}_{\mu}\right|^{l}Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (17)
to the molecular basis set. $G_{l,m,\sigma}^{\mu}$, these help make configuration integrals cheap. Here $\mathcal{N}_{l}$ is a normalization constant and cartesian Gaussians are also possible. This is a basis set for many small molecules. Furthermore Gaussians Coulomb integrals may be efficiently computed [28, 18] so there are “few” coulomb integrals involving complicated functions [29]. In particular if we consider the number of LO or lo orbitals as order one this becomes a three center problem at worst. We note that contracted Gaussian Functions (CGF) are also possible [27, 18]
III.2 Slo and SLO basis sets
We now consider a second hybrid basis set made of Slater type orbitals and localized orbitals. We first consider basis functions of the form given in Eqs. (14) or (16). We augment them with basis wave functions of the form:
$S_{l,m,\xi}^{\mu}=\frac{1}{\mathcal{N}_{l}}\exp\left(-\xi\left|\mathbf{r-r}_{\mu}\right|\right)\left|\mathbf{r-r}_{\mu}\right|^{l}Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\mu}}\right)$ (18)
While four center integrals are very difficult with this basis set three center integrals are numerically manageable [27]. Below we describe practical methods to achieve accurate three center decompositions.
III.2.1 Practical methods for implementation(three center decompositions)
We consider standard shape functions $\omega_{\kappa}$ for the density and add to them the following additional shape functions:
$\omega_{\mu Elm}^{E^{\prime}l^{\prime}m^{\prime}}\left(\mathbf{r}\right)=\Phi_{lo/LO}^{\mu E^{\prime}l^{\prime}m^{\prime}*}\left(\mathbf{r}\right)\cdot\Phi_{lo/LO}^{\mu Elm}\left(\mathbf{r}\right)$ (19)
this allows for hybridization between the atomic orbitals such as $sp^{3}$ hybridization common to Carbon (if this effect is not important for the molecule just the diagonal terms suffice). Furthermore for notational reasons we absorb the label $\mu;Elm;E^{\prime}l^{\prime}m^{\prime}$ into $\kappa$. We introduce the projectors $P_{\lambda}$ such that
$\int d^{3}\mathbf{r}P_{\lambda}^{*}\left(\mathbf{r}\right)\omega_{\kappa}\left(\mathbf{r}\right)=\delta_{\lambda\kappa}$ (20)
Where for example we may choose:
$M_{\alpha\beta}$ $=\int d^{3}\mathbf{r}\omega_{\alpha}^{*}\left(\mathbf{r}\right)\omega_{\beta}\left(\mathbf{r}\right)$
$P_{\lambda}\left(\mathbf{r}\right)$ $=\sum_{\kappa}\left[M^{-1}\right]_{\lambda\kappa}\omega_{\kappa}\left(\mathbf{r}\right)$ (21)
Then if we introduce the single particle density $\rho\left(\mathbf{r}\right)$ we may write that [27]
$\rho\left(\mathbf{r}\right)$ $\cong\sum_{\kappa}\left[\int d^{3}\mathbf{r}^{\prime}\rho\left(\mathbf{r}^{\prime}\right)P_{\kappa}^{*}\left(\mathbf{r}^{\prime}\right)\right]\omega_{\kappa}\left(\mathbf{r}\right)\equiv\sum_{\kappa}c_{\kappa}\omega_{\kappa}\left(\mathbf{r}\right)$
$c_{\kappa}$ $=\int d^{3}\mathbf{r}^{\prime}\rho\left(\mathbf{r}^{\prime}\right)P_{\kappa}\left(\mathbf{r}^{\prime}\right)$ (22)
We now introduce more notation for the single particle basis set where $\left\{\nu\right\}=\left\{Elm\right\}\cup\left\{\mu,l,m,\xi\right\}$ and write the eigenstates are given by
$\left|\lambda_{n}\right\rangle=\sum_{\nu}\lambda_{n\nu}\left|\psi_{\nu}\right\rangle$ (23)
We may now introduce [27]:
$P_{\mu\nu}=\sum_{n}f\left(\varepsilon_{n}\right)\lambda_{n\mu}^{*}\lambda_{n\nu}$ (24)
Then the density is given by:
$\rho\left(\mathbf{r}\right)$ $=\sum_{\mu\nu}P_{\mu\nu}\psi_{\mu}^{*}\left(\mathbf{r}\right)\psi_{\nu}\left(\mathbf{r}\right)$
$\rho\left(\mathbf{r}\right)$ $\cong\sum_{\kappa}\sum_{\mu\nu}P_{\mu\nu}\left[\int d^{3}\mathbf{r}^{\prime}\psi_{\mu}^{*}\left(\mathbf{r}^{\prime}\right)\psi_{\nu}\left(\mathbf{r}^{\prime}\right)P_{\kappa}^{*}\left(\mathbf{r}^{\prime}\right)\right]\omega_{\kappa}\left(\mathbf{r}\right)$ (25)
Which is a three center integral. Furthermore the KS equation is now given by:
$H_{\alpha\beta}^{KS}$ $=\left\langle\psi_{\alpha}\mid\frac{-\nabla^{2}}{2m}-\sum_{\mu}\frac{Z_{\mu}e^{2}}{\left|\mathbf{R}_{\mu}-\mathbf{r}\right|}+V_{XC}\left(\mathbf{r}\right)\mid\psi_{\beta}\right\rangle$
$+\sum_{\kappa}c_{\kappa}\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\omega_{\kappa}\left(\mathbf{r}_{1}\right)\frac{e^{2}}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}\psi_{\alpha}^{*}\left(\mathbf{r}_{2}\right)\psi_{\beta}\left(\mathbf{r}_{2}\right)$ (26)
Which is another three center integral. As such it is possible to do practical calculations for small molecules within DFT for SLO and Slo basis sets.
III.3 Plo and PLO basis sets for crystalline solids
@@
Here P in Plo and PLO stands for plane wave while lo and LO stands for localized orbitals. We first consider basis functions of the form given in Eqs. (14) or (16) . We also add plane wave
@@
to the solid state basis set. This is a basis set for many crystalline solids. However there are many competitor basis sets [9, 11, 12, 13, 14, 17, 3, 8, 7, 5, 6, 1, 10, 15, 16].
@@ IV Conclusions
In this work - inspired by the fact that many pseudopotential methods are all electron methods in disguise with slowly oscillating basis wave functions and special basis wave functions to account for the environment of the nucleus - we proposed new basis sets for small molecules and crystalline solids. In this basis set we use the LO or lo wave functions introduced recently in [26] (although regular LO or lo basis wave functions will also do [5, 6, 1]) combined either with Gaussians (GTOs and CGFs) and slater basis sets - STO - (for molecules) or plane waves (for solids) to obtain a total basis for the system. This would allow for initial calculations for many molecules as the four center calculation is done for Gaussians and two center density density is done for very generic functions [2, 29, 18] and for slater type orbitals it is possible to introduce appropriate density shape functions. This should help open small molecules for theoretical DFT or HF like calculations and explorations.